Each of the four independent situations below describes a sales-type lease in which annual lease payments of $145,000 are payable at the beginning of each year. Each is a finance lease for the lessee. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)Situation 4 13% 12% 7 12% Lease term (years) Lessors and lessees interest rate Residual value: Estimated fair value GuarantTABLE 1 Future Value of $1 FV = $1 (1 + i) n/i 1.0% 1.5% 2.0% 2.5% 3.0% 3.5% 4.0% 4.5% 5.0% 5.5% 1 1.01000 1.01500 1.02000 1.TABLE 2 Present Value of $1 PV 0.0% 0.91743 0.0% 0.90909 11.0% 0.90090 12.0% 0.89286 20.0% 0.83333 n/i 1.0% 1.5% 2.0% 2.5% 3.TABLE 3 Future Value of an Ordinary Annuity of $1 FVA - ^1 +in-1 n/i 1.0% 1.5% 2.0% 2.5% 3.0% 3.5% 4.0% 1 1.0000 1.0000 1.000TABLE4 Present Value of an Ordinary Annuity of $1 PVA -(1 + ir PVA ni 1.0% 1 0.99010 2 1.97040 3 2.94099 4 3.90197 5 4.85343TABLE 5 Future Value of an Annuity Due of $1 FVAD - (1 + i) - 1 (1 + i) n/i 1 2 3 4 5 1.0% 1.0100 2.0301 3.0604 4.1010 5.152TABLE 6 Present Value of an Annuity Due of $1 PVAD ---+*P}x^1 + ni 1 2 3 4 1.0% 1.00000 1.99010 2.97040 3.94099 1.5% 1.00000
Answer #1

Situations

1

2

3

4

A

The Lessor’s:

1

Lease payments

870000

870000

1023900

1084000

2

Gross investment in the lease

870000

929000

1032800

1143000

3

Net investment in the lease

680906

683337

749206

799055

B

The lessee’s:

4

Lease payments

870000

870000

1023900

1084000

5

Right-of-use asset

680906

654998

745180

772367

6

Lease payable

680906

654998

745180

772367

Situation 1

Lease payments = (145000*6) = 870000

Gross investment in the lease = (145000*6) = 870000

Net investment in the lease = 145000*4.69590= 680906

Right-of-use asset = 145000*4.69590= 680906

Lease payable = 145000*4.69590= 680906

Present value annuity due factor of $1 @ n= 6 , i=11% = 4.69590

Situation 2

Lease payments = (145000*6) = 870000

Gross investment in the lease = (145000*6) + 59000 = 929000

Net investment in the lease = (145000*4.51723)+ (59000*0.48032) = 683337

Right-of-use asset = 145000*4.51723= 654998

Lease payable = 145000*4.51723= 654998

Present value annuity due factor of $1 @ n= 6 , i=13% = 4.51723

Present value factor of $1 @ n= 6 , i=13% = 0.48032

Situation 3

Lease payments = (145000*7)+8900 = 1023900

Gross investment in the lease = (145000*7)+8900+8900 = 1032800

Net investment in the lease = (145000*5.11141)+(17800*0.45235)= 749206

Right-of-use asset = (145000*5.11141)+(8900*0.45235)= 745180

Lease payable = (145000*5.11141)+(8900*0.45235) = 745180

Present value annuity due factor of $1 @ n= 7 , i=12% = 5.11141

Present value factor of $1 @ n= 7 , i=12% = 0.45235

Situation 4

Lease payments = (145000*7)+69000 = 1084000

Gross investment in the lease = (145000*7)+69000 +59000= 1143000

Net investment in the lease = (145000*5.11141)+(128000*0.45235)= 799055

Right-of-use asset = (145000*5.11141)+(69000*0.45235)= 772367

Lease payable = (145000*5.11141)+(69000*0.45235) = 772367

Present value annuity due factor of $1 @ n= 7 , i=12% = 5.11141

Present value factor of $1 @ n= 7 , i=12% = 0.45235